### v1

@@ 4.3.1 Methodology
In RQ3, our goal is to understand how PTMs collaborate with other learned components, i.e., scratch-trained models or additional PTMs, to form learned behaviors within ML-enabled systems. Similar to RQ2, we first conduct a literature review to identify commonly discussed model interaction patterns across deep learning subfields, which forms a theoretical basis for our analysis. Next, we manually examine the 401 target projects and identify 200 that include more than one model; these form the dataset for RQ3. For each selected project, we analyze whether any PTM interacts with other model(s) within its reuse pipeline. We then map the observed interaction patterns to those identified in the literature and extend the taxonomy as novel patterns emerge.
@@
To collect relevant work, we focus on survey papers because they consolidate canonical design patterns and offer generalized architectural taxonomies, unlike individual empirical studies. Here, we only use Google Scholar because we prioritize popular architectural designs over whether they are published in SE or ML venues. Similar to RQ2, we use the Publish or Perish tool to query Google Scholar with the following search string:
@@ 4.3.2 Result
Interaction Types in PTM Reuse Pipeline. Table 5 summarizes the four interaction types identified in the 200 target projects. Feature Handoff and Feedback Guidance are the most common, each accounting for more than 40% of observed examples, while Evaluation and Post-Processing are less frequent. This indicates that PTM reuse in OSS pipelines is dominated by patterns involving either direct feature passing or bidirectional feedback during training. Compared with prior literature, which covers two interaction types without focusing on PTMs, our study identifies four interaction types, two more than previously recognized, and highlights finer-grained cross-stage PTM-specific patterns. Next, we describe the characteristics and provide examples for each category.
@@
In this interaction type, a PTM’s output is employed to assess the quality of another model’s outputs. These interactions typically occur post hoc (at the evaluation stage). The PTM provides objective metrics or diagnostic insights but does not influence model behavior directly. This pattern is common in generative systems. For instance, Inception-v3 and CLIP are often used to compute Fréchet Inception Distance (FID) or CLIP-based similarity scores for evaluating images or text-to-image models. It also appears in NLP applications. For instance, pretrained language models (e.g., GPT) can be used to generate paraphrased or alternative entity descriptions, which are then fed into a named entity recognition (NER) system to test its robustness. These models are not involved in the training or adaptation of the NER model but serve only as tools to generate diverse evaluation inputs. [^7] For example, see https://github.com/IBM/zshot Figure 11 illustrates a typical setup.
@@ 5 Discussion and Implications
This section reflects on our study’s findings from three perspectives. First, we discuss the broader conceptual implications for software dependencies, highlighting the paradigm shift from Software Dependencies 1.0 to Software Dependencies 2.0. Next, we examine the maintenance challenges that arise from PTM dependencies, pipeline complexity, and multi-model interactions in ML-enabled systems. Finally, we outline actionable implications for key ML stakeholders, including researchers, developers, PTM hub providers, and MLOps platform providers.
@@
Ease of Integration Simple; typically an init() function or small configuration file (Decan et al., 2019). High configuration cost; involves code API + model artifacts (§4.1), multiple parameters, adapter/head selection (§4.2), and multi-model coordination (§4.3).
@@
Risk & Vulnerability Code-level: CVEs, API breaks, version conflicts (Decan et al., 2019; Sammak et al., 2023). Model-level: contamination, adversarial risk, unsafe fine-tuning, license conflicts, attack transferability (Vassilev, 2024; Gao et al., 2025).
@@
Our analysis highlights a fundamental, qualitative shift between conventional software dependencies (Software Dependencies 1.0) and the paradigm of model-centric dependencies (Software Dependencies 2.0). Our measurements echo another recent work by  Banyongrakkul et al. (2025), which emphasizes that integrating PTMs into downstream applications introduces unique challenges. We believe these challenges arise from a shift from code-centric to model-centric artifacts, motivating our introduction of the term Software Dependencies 2.0. Table 7 summarizes these distinctions based on our findings. While Software Dependencies 1.0 are grounded in code-centric artifacts, the reuse of PTMs in ML-enabled systems introduces model-centric components whose behavior is shaped by the learned PTM, their reuse pipelines and interactions with other models. This evolution redefines the unit of dependency management from a simple library + version to a more complex composite: model + provenance + reuse pipeline + interaction.
@@
Ease of Integration
In Software Dependencies 1.0, code libraries often require minimal setup, typically a simple init() function or a small configuration file. In contrast, Software Dependencies 2.0 involve both code APIs and model artifacts (weights, configurations, adapters, etc.). PTM reuse requires managing multiple parameters, specifying model names with versions (Table 7), selecting heads or adapters, adjusting configurations (Table 4), and coordinating multi-model pipelines (Table 6). This configuration complexity highlights that PTM dependencies are far more complex to integrate than code-centric dependencies.
@@
Beyond Code Vulnerabilities: Model-Specific Risks
This one is our synthesis of other literature, but does not show up in our data. Software Dependencies 1.0 focus on code-level risks such as CVEs, API breaks, and version conflicts. In contrast, Software Dependencies 2.0 introduce model-specific vulnerabilities, such as data contamination, adversarial susceptibility, unsafe fine-tuning, license conflicts (Decan et al., 2019; Sammak et al., 2023), and attack transferability. These risks propagate not only via code but also through reuse pipelines and interaction edges, requiring security analysis that extends to models and their provenance.
@@ 5.2 PTM-Centric Maintenance Challenges in ML-Enabled Systems
While prior work has examined technical debt in ML systems (Sculley et al., 2015; Bogner et al., 2021; Gesi et al., 2022; Recupito et al., 2024), our study highlights new challenges that arise from the shift from training models from scratch to reusing PTMs or combining PTMs with scratch-trained models. In Table 8, we map the challenges categorized in our study to those reported in prior research on technical debt and other engineering challenges in ML-enabled systems. We then elaborate on each challenge category, linking it to our empirical findings.
@@
Challenge Category Conventional ML (Prior Work) PTM Reuse (This Study)
@@
Accurate and consistent PTM referencing is essential for effective dependency tracking, especially given that 52.6% of projects in our dataset utilize multiple PTMs rather than a single one. In practice, however, PTM references are often scattered across source code, documentation, and configuration files, with only 21.2% of projects declaring all PTMs outside the codebase (§4.1). Additionally, only 12% of projects specify the version of PTM. This decentralized and informal linking with missing version identifiers results in incomplete provenance metadata, creating dependency debt that complicates the long-term management of model dependencies. PTM reuse also suffers from configuration debt, i.e., the uncontrolled scattering of model metadata across multiple scripts, configuration files, and code blocks (§4.1), and from reproducibility debt, as downstream behavior can change unexpectedly when upstream models are updated due to the absence of structured metadata such as model names and versions. These findings align with prior work on undeclared consumer dependencies in ML systems (Sculley et al., 2015)
@@ 5.3 Implications
Another important direction is extending software quality frameworks (Gesi et al., 2022; Studer et al., 2021) to account for the unique PTM lifecycle. Attributes such as fairness, interpretability, and robustness can evolve as PTMs are reused in pipelines (Bender et al., 2021). Systematic measurement and management of these qualities will enable new evaluation methodologies and engineering practices. Additionally, Software Dependencies 2.0 (§4.1) highlights future research opportunities in managing interchangeable and complementary PTM dependencies, including safe model substitution, coordination across heterogeneous models, and designing frameworks that support both substitution and composition in complex pipelines.
Finally, our results suggest complementary opportunities for ML research, including methods for fine-tuning with clear provenance, model-editing strategies, and metrics for multimodel pipeline evaluation. Together, these directions provide an empirically grounded agenda for both SE and ML research communities, promoting more sustainable, reproducible, and trustworthy reuse of pretrained models in software systems. We provide a foundational base, while future work should investigate data–model–code entanglement across cross-stage model interactions.
Future work should also examine PTMs accessed via model-centric platforms (MCPs) (Krishnan, 2025), agent hubs (Ehtesham et al., 2025), and commercial APIs (OpenAI, 2023b). These sources may exhibit different dependency and reuse patterns compared to open-source hubs, and analyzing them could help generalize our findings across broader ML software ecosystems.
@@
Our analysis highlights maintenance challenges that arise when PTMs are integrated into projects. Developers should adopt structured practices to manage this integration. For dependency tracking and versioning (§4.1), models should be consistently declared, version-pinned, and treated as evolving software components. Across pipeline stages (§4.2), adaptations such as task-specific head additions or layer modifications should be documented, and automated tools could help generate such documentation (e.g., visualizations of model changes, integration points, and interfaces with other components) based on PTM reuse patterns and source code. Integration tests should be included in CI/CD pipelines to ensure reliability and consistency. For multi-model interactions (§4.3), outputs should be monitored to prevent downstream failures. Standardized tooling for dependency management, pipeline testing, and performance monitoring can further support reproducibility and stability in PTM-based systems.
@@
Our study highlights gaps in managing Software Dependencies 2.0. Traditional package managers and dependency graphs are insufficient for PTM reuse, and existing PTM “managers”, such as Hugging Face, offer only partial support: they lack standardized manifest files, reproducible builds, and complete visibility into model versions or third-party library dependencies, leaving downstream projects with incomplete dependency information. To help dependency management, (§4.1), hubs can provide immutable, versioned snapshots, structured changelogs, and rich metadata capturing model provenance. RQ2 (§4.2) showed diverse adaptation stages across three pipeline types, emphasizing that PTM reuse is rarely plug-and-play. Hubs can support developers by offering tools to visualize or compare model differences, highlight adaptation changes, and track compatibility across versions, and allow downloads of model components. Integration with CI/CD can help with automated testing of adapted models, dependency checks, and safe deployment of PTM updates. Finally, RQ3 (§4.3) demonstrated frequent cross-stage interactions. Hubs can provide metadata or APIs exposing interaction patterns, helping developers anticipate and manage potential incompatibilities in complex, multi-model pipelines. Together, these observations highlight that dedicated tooling for PTM dependency management should be developed.
@@
MLOps systems must become PTM-aware, treating the pipeline itself as a first-class component, extracting, tracking, and validating model metadata during execution—including version, adaptation and fine-tuning history. To manage pipeline complexity (§4.2), platforms can provide metrics for interaction quality, such as reuse intensity, adaptation overlap, and dependency distance. Observability mechanisms help track and validate PTMs within complex pipelines, making it easier to detect issues when models interact or when downstream data depends on upstream outputs (Pineau et al., 2021). Integrating these checks into CI/CD workflows will allow automatic downstream validation whenever a PTM is swapped or updated, supporting more robust and maintainable multi-PTM pipelines and cross-stage interactions (§4.3). Such automated workflow enforcement resonates with challenges reported in Workflows as Code systems, where defining, executing, and monitoring complex pipelines also require systematic observability and validation practices (Yasmin et al., 2024).
@@ 6.3 External Validity
We chose PeaTMOSS because it provides a curated, large-scale, and publicly accessible dataset, enabling reproducible analysis. Since PTM technology and usage patterns evolve rapidly—with advances in architecture, training methods, and dependency practices—our findings may become outdated. However, because our study focuses on general patterns of PTM reuse, it is not domain-specific and captures design principles applicable across diverse software projects, making the results broadly generalizable. Future studies should revisit these trends over time and extend the analysis to newer projects, additional domains, and different stages of software development to assess stability and generality.

### v2

@@ 4.3.1 Methodology
In RQ3, our goal is to understand how PTMs collaborate with other learned components, i.e., scratch-trained models or additional PTMs, to form learned behaviors within ML-enabled systems.
Similar to RQ2, we first conduct a literature review to identify commonly discussed model interaction patterns across deep learning subfields, which forms a theoretical basis for our analysis. Next, we manually examine the 401 target projects and identify 200 that include more than one model; these form the dataset for RQ3. For each selected project, we analyze whether any PTM interacts with other model(s) within its reuse pipeline. We then map the observed interaction patterns to those identified in the literature and extend the taxonomy as novel patterns emerge.
@@
To collect relevant work, we focused on survey papers, as they summarize recurring design patterns and provide generalized taxonomies, rather than reporting isolated empirical findings. For this step, we relied exclusively on Google Scholar to identify highly cited studies on PTM workflow architectures and interaction patterns across diverse venues (e.g., software engineering, machine learning, and domain-specific conferences). Compared to venue-specific digital libraries, Google Scholar offers broader cross-disciplinary coverage and citation-based ranking, which aligns with our goal of capturing widely adopted reusable patterns. Although this strategy may miss some relevant work, we believe it is sufficient to cover commonly recognized patterns. Similar to RQ2, we use the Publish or Perish tool to query Google Scholar with the following search string:
@@
We selected keywords based on deep learning terminology (deep learning, large language model, foundation model, pre-trained model) and interaction-related terms (interaction, multi-modal, multi-model, architecture, fusion). We focused on survey papers because they summarize recurring design patterns and provide generalized architectural taxonomies, rather than reporting isolated empirical findings. During scoping, we tested variations and found the final query balanced coverage and relevance well.
@@ 4.3.2 Result
Interaction Types in PTM Reuse Pipeline. Table 5 summarizes the four interaction types identified in the 200 target projects. Feature Handoff and Feedback Guidance are the most common, each accounting for more than 40% of observed examples, while Evaluation and Post-Processing are less frequent. This indicates that PTM reuse in OSS pipelines is dominated by patterns involving either direct feature passing or bidirectional feedback during training. Compared with prior literature, which covers two interaction types without focusing on PTMs, our study identifies four interaction types, two more than previously recognized, and highlights finer-grained cross-stage PTM-specific patterns.
Next, we describe the characteristics and give examples for each category.
@@
In this interaction type, a PTM’s output is employed to assess the quality of another model’s outputs. These interactions typically occur post hoc (at the evaluation stage). The PTM provides objective metrics or diagnostic insights but does not influence model behavior directly. This pattern is common in generative systems. For instance, Inception-v3 and CLIP are often used to compute Fréchet Inception Distance (FID) or CLIP-based similarity scores for evaluating images or text-to-image models. It also appears in NLP applications. For instance, pretrained language models (e.g., GPT) can be used to generate paraphrased or alternative entity descriptions, which are then fed into a named entity recognition (NER) system to test its robustness. These models are not involved in the training or adaptation of the NER model but serve only as tools to generate diverse evaluation inputs. [^8] For example, see https://github.com/IBM/zshot. Figure 12 illustrates a typical setup.
@@ 5 Discussion and Implications
This section reflects on our study’s findings from three perspectives. First, we discuss the broader conceptual implications for software dependencies, focusing on the distinctions between Software Dependencies 1.0 and Software Dependencies 2.0. Next, we examine the maintenance challenges that arise from PTM dependencies, pipeline complexity, and multi-model interactions in ML-enabled systems. Finally, we outline actionable implications for key ML stakeholders, including researchers, developers, PTM hub providers, and MLOps platform providers.
@@
Integration Often simple; typically an init() function or small configuration file (Decan et al., 2019) Often involves high configuration cost; involves code API + model artifacts (§4.1), multiple parameters, adapter/head selection (§4.2), and multi-model coordination (§4.3).
@@
Our analysis highlights notable differences between conventional software dependencies (Software Dependencies 1.0) and the paradigm of model-centric dependencies (Software Dependencies 2.0). Our measurements echo another recent work by  Banyongrakkul et al. (2025), which emphasizes that integrating PTMs into downstream applications introduces unique challenges. We believe these challenges arise from a shift from code-centric to model-centric artifacts, motivating our introduction of the term Software Dependencies 2.0. Table 7 summarizes these distinctions based on our findings. While Software Dependencies 1.0 are grounded in code-centric artifacts, the reuse of PTMs in ML-enabled systems introduces model-centric components whose behavior is shaped by the learned PTM, their reuse pipelines and interactions with other models. This evolution extends the unit of dependency management from a simple library + version to a more complex composite: model + provenance + reuse pipeline + interaction.
@@
Integration
In Software Dependencies 1.0, code libraries often require minimal setup, typically a simple init() function or a small configuration file. In contrast, Software Dependencies 2.0 involve both code APIs and model artifacts (weights, configurations, adapters, etc.). PTM reuse requires managing multiple parameters, specifying model names with versions (Table 8), selecting heads or adapters, adjusting configurations (Table 4), and coordinating multi-model pipelines (Table 6). This configuration complexity suggests that PTM dependencies are typically more complex to integrate than code-centric dependencies.
@@
Debate: Does This Concept Merit the Label “Software Dependencies 2.0”?
Our use of the term “Software Dependencies 2.0” is intended as an analytical abstraction rather than a categorical replacement of conventional dependency models. We do not claim that ML-enabled systems discard all properties of Software Dependencies 1.0, nor that every PTM integration exhibits all of the characteristics summarized in Table 7. Instead, our conceptualization captures a recurring pattern observed across our empirical analysis: that model-centric artifacts introduce semantic dependency edges, pipeline-embedded adaptation, cross-model interaction, and decentralized traceability that are not fully captured by conventional manifest-based dependency models. We intend “2.0” to reflect a substantive extension of the unit of dependency management, from “library + version” to “model + provenance + version + reuse pipeline + interaction”. Our extension is grounded in the structural regularities we observed across 401 projects.
We acknowledge that a reader might examine any of the rows in Table 7 and point to counterexamples in conventional software. For example, not all conventional dependencies are simple to incorporate, probabilistic algorithms are well known to introduce nondeterminism, and plugin ecosystems can exhibit complex inter-package interactions. Our position follows the adage, “The exception proves the rule.” While individual 1.0 systems may display isolated instances of these properties, our findings indicate that such characteristics are structural and pervasive in PTM reuse. We assert not that PTM reuse introduces entirely novel phenomena, but that it systematically combines behavioral, architectural, and provenance complexity in ways that make existing dependency abstractions incomplete. We therefore offer the term Software Dependencies 2.0 as a conceptual tool to reason about this emerging pattern, and its merit ultimately rests on whether it clarifies analysis and supports more reliable engineering practice.
@@ 5.2 PTM-Centric Maintenance Challenges in ML-Enabled Systems
While prior work has examined technical debt in ML systems (Sculley et al., 2015; Bogner et al., 2021; Gesi et al., 2022; Recupito et al., 2024), our study suggests new maintenance challenges that arise from the shift from training models from scratch to reusing PTMs or combining PTMs with scratch-trained models. Recent PTM-specific studies have cataloged practitioner challenges from different angles: Tan et al. (2024) focus on PTM usage issues such as model selection, testing, while Banyongrakkul et al. (2025) classify developer-reported challenges across seven themes, including model I/O handling, dependency incompatibility, hardware constraints, and missing model resources. These works characterize what challenges developers encounter when working with PTMs. Our work complements theirs by examining how these challenges manifest as maintenance concerns within reuse pipelines: specifically, through dependency structures and documentation practices (RQ1), pipeline-stage organization and adaptation patterns (RQ2), and cross-model interactions (RQ3).
In Table 8, we map the challenges categorized in our study to those reported in prior research on technical debt and other engineering challenges in ML-enabled systems. We confirm maintenance-related aspects previously identified in conventional ML systems (e.g., configuration complexity, dependency management) while extending them with PTM-specific and pipeline-centric insights from RQs. Notably, several maintenance challenges are specific to the PTM reuse context and have not been reported in prior work, including model-data entanglement and interaction debt from cross-stage multi-model coordination. We then elaborate on each maintenance challenge category, linking it to our empirical findings.
@@
Maintenance Challenge Category Conventional ML (Prior Work) PTM Reuse (This Study)
@@
Accurate and consistent PTM referencing is essential for effective dependency tracking, especially given that 52.6% of projects in our dataset utilize multiple PTMs rather than a single one. In practice, however, PTM references are often scattered across source code, documentation, and configuration files, with only 21.2% of projects declaring all PTMs outside of source code (i.e., a README or configuration file)(§4.1). Additionally, only 12% of projects specify the version of PTM. This decentralized and informal linking with missing version identifiers results in incomplete provenance metadata, creating dependency debt that complicates the long-term management of model dependencies. PTM reuse also suffers from configuration debt, i.e., the uncontrolled scattering of model metadata across multiple scripts, configuration files, and code blocks (§4.1), and from reproducibility debt, as downstream behavior can change unexpectedly when upstream models are updated due to the absence of structured metadata such as model names and versions. These findings align with prior work on undeclared consumer dependencies in ML systems (Sculley et al., 2015) However, we acknowledge that PTM versioning is not commonly reported as a critical practitioner challenge in existing taxonomies (Banyongrakkul et al., 2025; Tan et al., 2024), suggesting it may represent a latent maintenance risk rather than an immediate usability concern for many developers.
@@ 5.3 Implications
Another important direction is extending software quality frameworks (Gesi et al., 2022; Studer et al., 2021) to account for the unique PTM lifecycle. Attributes such as fairness, interpretability, and robustness can evolve as PTMs are reused in pipelines (Bender et al., 2021). Systematic measurement and management of these qualities will enable new evaluation methodologies and engineering practices. Additionally, Software Dependencies 2.0 highlights future research opportunities in managing interchangeable and complementary PTM dependencies, including safe model substitution, coordination across heterogeneous models, and designing frameworks that support both substitution and composition in complex pipelines.
Finally, our results suggest complementary opportunities for ML research, including methods for fine-tuning with clear provenance (Meiklejohn et al., 2025), model watermarking to track PTM lineage across reuse (Xian et al., 2024; Cohen et al., 2025), model-editing strategies, and metrics for multi-model pipeline evaluation.
Together, these directions provide an empirically grounded agenda for both SE and ML research communities, promoting more sustainable, reproducible, and trustworthy reuse of pretrained models in software systems. We provide a foundational base, while future work should investigate data–model–code entanglement across cross-stage model interactions.
Future work should also examine PTMs accessed via model-centric platforms (MCPs) (Krishnan, 2025), agent hubs (Ehtesham et al., 2025), and commercial APIs (OpenAI, 2023b). While PTM reuse provides a canonical instantiation of Software Dependencies 2.0, other model-centric artifacts, such as hosted foundation models, retrieval-augmented pipelines, adapter libraries, and agentic compositions, may exhibit similar structural properties. These environments may surface different dependency declaration practices, provenance mechanisms, and interaction patterns compared to open-source hubs. Analyzing such systems would help assess the generality of Software Dependencies 2.0 as an abstraction and refine its boundaries across broader ML software ecosystems.
@@
Our analysis highlights maintenance challenges that arise when PTMs are integrated into projects. Developers should adopt structured practices to manage this integration. Models should be consistently declared, version-pinned, and treated as evolving software components to support dependency tracking and versioning (§4.1). Adaptations such as task-specific head additions or layer modifications should be documented across pipeline stages, and automated tools could help generate such documentation (e.g., visualizations of model changes, integration points, and interfaces with other components) based on PTM reuse patterns and source code (§4.2). Integration tests should be included in CI/CD pipelines to ensure reliability and consistency. Outputs from multi-model interactions should be monitored to prevent downstream failures (§4.3). Standardized tooling for dependency management, pipeline testing, and performance monitoring can further support reproducibility and stability in PTM-based systems.
@@
Our study highlights gaps in managing Software Dependencies 2.0. Traditional package managers and dependency graphs are insufficient for PTM reuse, and existing PTM “managers”, such as Hugging Face, offer only partial support: they lack standardized manifest files, reproducible builds, and complete visibility into model versions or third-party library dependencies, leaving downstream projects with incomplete dependency information. To help dependency management, hubs can provide immutable, versioned snapshots, structured changelogs, and rich metadata capturing model provenance. RQ2 (§4.2) showed diverse adaptation stages across three pipeline types, emphasizing that PTM reuse is rarely plug-and-play. Hubs can support developers by offering tools to visualize or compare model differences, highlight adaptation changes, and track compatibility across versions, and allow downloads of model components. Integration with CI/CD can help with automated testing of adapted models, dependency checks, and safe deployment of PTM updates. Finally, RQ3 (§4.3) demonstrated frequent cross-stage interactions. Hubs can provide metadata or APIs exposing interaction patterns, helping developers anticipate and manage potential incompatibilities in complex, multi-model pipelines. Together, these observations highlight that dedicated tooling for PTM dependency management should be developed.
@@
MLOps systems must become PTM-aware, treating the pipeline itself as a first-class component, extracting, tracking, and validating model metadata during execution—including version, adaptation and fine-tuning history. To manage pipeline complexity (§4.2), platforms can provide metrics for interaction quality, such as reuse intensity, adaptation overlap, and dependency distance. Observability mechanisms help track and validate PTMs within complex pipelines, making it easier to detect issues when models interact or when downstream data depends on upstream outputs (Pineau et al., 2021). Integrating these checks into CI/CD workflows will allow automatic downstream validation whenever a PTM is swapped or updated, supporting more robust and maintainable multi-PTM pipelines and cross-stage interactions (§4.3). Such automated workflow enforcement resonates with challenges reported in Workflows as Code systems, where defining, executing, and monitoring complex pipelines also require systematic observability and validation practices (Yasmin et al., 2025).
@@ 6.3 External Validity
We chose PeaTMOSS because it provides a curated, large-scale, and publicly accessible dataset, enabling reproducible analysis. Since PTM technology and usage patterns evolve rapidly—with advances in architecture, training methods, and dependency practices—our findings may become outdated.
However, because our study focuses on general patterns of PTM reuse, it is not domain-specific and captures design principles applicable across diverse software projects, making the results broadly generalizable. Future studies should revisit these trends over time and extend the analysis to newer projects, additional domains, and different stages of software development to assess stability and generality.
6.4 Conclusion Validity
Conclusion validity concerns whether our observed relationships and interpretations are well supported by the data. For project categorizations and the investigation of RQs, our taxonomies were developed through systematic sampling and manual analysis, which may not capture the full diversity of PTM usage patterns and inherently involve subjective judgment. In particular, for RQ2 and RQ3, our systematic review process required manual inspection and categorization of code patterns to identify stages and interaction types, introducing potential researcher bias.
To mitigate these threats, we adopted several safeguards. First, we developed explicit coding guidelines with representative examples for each category. Second, we conducted inter-rater reliability checks, with multiple researchers independently coding subsets of the data. Third, we held regular meetings to discuss disagreements and resolve ambiguities. Fourth, we iteratively refined our taxonomies as new patterns emerged during analysis. When applying the taxonomies to larger samples, we further validated them through pilot applications and documented edge-case resolutions (§ 3.2, § 4.1, § 4.2, § 4.3). Despite these safeguards, subjective judgment remains inherent in qualitative analysis, and alternative interpretations of certain code patterns are possible.
Finally, our interaction taxonomy is grounded in patterns observed within our dataset and may not be exhaustive. For example, while we identified parallel feature handoff patterns, we did not observe ensemble-style mechanisms (e.g., voting or averaging) in our sample, which may reflect our dataset’s characteristics rather than the absence of such patterns in practice.
Our versioning analysis examined multiple sources beyond just model-loading statements, including configuration files (YAML/JSON), README files, and requirements specifications. However, we acknowledge that our analysis did not extend to all possible version specification channels such as Docker images, Conda environments, or CI/CD configurations.
Our categorization of multi-PTM relationships uses high-level categories (interchangeable, complementary); finer-grained subcategories based on architecture, modality, or functional role could reveal additional patterns, but were not pursued due to concerns about coding complexity and inter-rater reliability. Future work could explore finer-grained subcategories of complementary PTM relationships based on architecture, modality, and functional role to identify potential connections between these dimensions.
Our comparison between Software Dependencies 1.0 and 2.0 is based on empirical observations from open-source projects rather than direct practitioner validation. Some challenges, such as versioning, may represent latent risks rather than immediate practical concerns. A user study validating these findings from a practitioner perspective is an important direction for future work.